OpenStudy (anonymous):

let f(x)=2-5x+11x^2. Calculate f(a+h)

OpenStudy (anonymous):

\[f \left( a+h \right)=2-5\left( a+h \right)+11\left( a+h \right)^{2}\] solve and get the solution.